OpenStudy (anonymous):

Someone I need help..! If the force of 12 N is applied to an object with a mass of 2 kg, the object will accelerate at 0.17 m/s². 24 m/s². 6 m/s². 12 m/s².

OpenStudy (anonymous):

I think there's a formula for this just not sure.

OpenStudy (mathstudent55):

F = ma

OpenStudy (mathstudent55):

F = force, m = mass, a = acceleration

OpenStudy (anonymous):

okay so it would look like 12= 12 something? I dont know what that is yet.

OpenStudy (mathstudent55):

F = ma, you have the force and the mass and are looking for the acceleration, so solve F = ma for a. Then plug in the values you have and get a value for a.

OpenStudy (anonymous):

Well I kindof understand that but not all the way. I don't understand how to plug in the numbers to the formula. I'm not good at formulas.

OpenStudy (mathstudent55):

a = F/m

OpenStudy (mathstudent55):

a = F/m = 12 N/2 kg = 6 (kg-m/s^2)/kg = 6 m/s^2

OpenStudy (anonymous):

Oh lord, thats a lot of working out...

OpenStudy (anonymous):

OHHH so is the answer 6???

OpenStudy (raden_zaikaria):